OpenStudy (anonymous):

If x=4tanθ, use trigonometric substitution to write √(16+x^2) as a trigonometric function of θ, where 0<θ<π/2

OpenStudy (anonymous):

yes, first substitute \[4\tan \theta\] for x

OpenStudy (anonymous):

so you have\[\sqrt{16+16\tan ^{2}\theta}\]

OpenStudy (anonymous):

I am assuming that you would plug in 4tanθ to get: ` √(16+(4tanθ)^2) just not sure where to go from there

OpenStudy (anonymous):

you can factor out a 16 and take it out of the radical \[\sqrt{16(1+\tan ^{2}\theta)} = 4\sqrt{(1+\tan ^{2}\theta)}\]

OpenStudy (anonymous):

now you need to make a trig substitution. \[1+\tan ^{2}\theta=\sec ^{2}\theta\]

OpenStudy (anonymous):

OK... so when you say take it out of the radical you mean...? sorry this is where I get confused

OpenStudy (anonymous):

since we had \[\sqrt{16(1+\tan ^{2}\theta)}\] we can evaluate the square root of 16

OpenStudy (anonymous):

Ohhh ok

OpenStudy (anonymous):

and take it out of the square root

OpenStudy (anonymous):

so then the final step would be \[4\sqrt{\sec ^{2}\theta}=4\sec \theta\]

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

sort of... I'm trying to figure out how you got 1+tan^2θ=sec^2θ to turn into 4√(sec^2θ)=4secθ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[4\sqrt{1+\tan ^{2}\theta}=4\sqrt{\sec ^{2}\theta}\]

OpenStudy (anonymous):

I first just replaced 1+tan^2θ with sec^2θ

OpenStudy (anonymous):

then, we can take the square root of sec^2θ

OpenStudy (anonymous):

which is just secθ

OpenStudy (anonymous):

oh ok that makes sense thanks!

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

say I had one with cot.... √(9+(3cot^2θ) √(9+9cotθ)?

OpenStudy (anonymous):

or I mean √(9+9cot^2θ)?

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

ok then I think....1√(1+cot^2θ) how do I know what to substitute then?

OpenStudy (anonymous):

Or I mean 9√(1+cot^2θ)